# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ D.1 Market-Clearing
Let $Z_{n}(\bm{w},p)=\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(w_{i},\bm{p})$ and $P_{n}(\bm{w})=\arg\min\limits_{p}|Z_{n}(\bm{w},p)|$ for $\bm{w}\in\{0,1\}^{n}$. In cases where there are multiple minimizers, choose one by some deterministic rule.
@@
is non-empty with probability $1-e^{-cn}$. And $P_{n}(\bm{w})\in\mathcal{S}_{\mathbf{w}}$ when it is non-empty.
@@
We have now verified that $\mathcal{S}_{\mathbf{w}}$ is non-empty with probability $1-e^{-cn}$, for some $c>0$, which completes the proof.
D.2 Convergence Rate of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$
@@
D.3 Observing a Sub-Sample of the Market
@@
D.4 Alternative Estimator of AIE Variance
In analyzing the asymptotic properties of $\sigma^{2}_{I}$ in the proof of Theorem 7 in Appendix B.5, we dropped a term that was asymptotically negligible, which was $\sqrt{n}h_{n}({\gamma}^{\top}(\hat{\tau}^{z}_{\text{ADE}}-\tau^{*,z}_{\text{ADE}})).$
$\tilde{\sigma}^{2}_{I}$ adds a plug-in estimator for the second term, including an estimator for the variance of $\hat{\tau}^{z}_{\text{ADE}}$. $\tilde{\sigma}^{2}_{I}=\hat{\sigma}^{2}_{I}+h^{2}_{n}\hat{\gamma}^{\top}\hat{\sigma}^{2}_{z,D}\hat{\bm{\gamma}},$ where $\hat{\sigma}^{2}_{z,D}=\frac{1}{n}\sum\limits_{i=1}^{n}B_{i}B_{i}^{\prime}$ and the $J\times 1$ vector $\hat{B}_{i}=\frac{W_{i}\hat{\varepsilon}^{z}_{i}(1)}{\hat{\pi}}-\frac{(1-W_{i})\hat{\varepsilon}^{z}_{i}(0)}{1-\hat{\pi}}-[\hat{\xi}_{z1}-\hat{\xi}_{z0}]^{\top}\hat{\xi}_{z}^{-1}{Z}_{i}$. ${\varepsilon}^{z}_{i}(1)=Z_{i}-\frac{1}{n_{w}}\sum\limits_{i:W_{i}=w}Z_{i}$. Last, the $J\times J$ matrix $\hat{\xi}_{zw}$ for $w\in\{0,1\}$ are estimated from regressions of $Z_{i}$ on $U_{i}$ using only observations such that $W_{i}=w$. We found in simulations that this second-order correction led to better coverage properties at smaller sample sizes.

### v4

@@ D.1 Market-Clearing
Let $Z_{n}(\bm{w},p)=\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(w_{i},p)$ and $P_{n}(\bm{w})=\arg\min\limits_{p}|Z_{n}(\bm{w},p)|$ for $\bm{w}\in\{0,1\}^{n}$. In cases where there are multiple minimizers, choose one by some deterministic rule.
@@
is non-empty with probability $1-e^{-c_{1}n}$. And $P_{n}(\bm{w})\in\mathcal{S}_{\mathbf{w}}$ when it is non-empty.
@@
We have now verified that $\mathcal{S}_{\mathbf{w}}$ is non-empty with probability $1-e^{-c_{1}n}$, for some $c_{1}>0$, which completes the proof.
D.2 Tighter Conservative Bound for $\bar{\sigma}^{2}_{D}$
First, introduce the notation $a_{i}(1)=\epsilon_{i}(1)-\pi\Delta_{i}(1,p^{*}_{\pi})$, and $a_{i}(0)=\epsilon_{i}(0)+(1-\pi)\Delta_{i}(0,p^{*}_{\pi}).$ Under the augmented randomized experiment, an estimator is available for the following variance.
$\tilde{\sigma}^{2}_{D}=\mathbb{E}\left[\frac{(1-\pi)}{\pi}a_{i}^{2}(1)+\frac{\pi}{1-\pi}a_{i}^{2}(0)+2\mathbb{E}[a_{i}(1)^{2}]^{1/2}\mathbb{E}[a_{i}(0)^{2}]^{1/2}\right].$
In an extension of the classical results in Neyman (1923), and described in detail in Section 2.1 of Aronow et al. (2014), we can show that this variance $\tilde{\sigma}^{2}_{D}$ is conservative for $\bar{\sigma}^{2}_{D}$ and is tighter than $\sigma^{2}_{D}$.
$\begin{split}\sigma^{2}_{D}&=\mathbb{E}\left[\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},p^{*}_{\pi})\right)^{2}\right],\\ &=\mathbb{E}\left[\left(\frac{W_{i}a_{i}(1)}{\pi}-\frac{(1-W_{i})a_{i}(0)}{1-\pi}\right)^{2}\right],\\ &=\mathbb{E}\left[\frac{a_{i}(1)^{2}}{\pi}\right]+\mathbb{E}\left[\frac{a_{i}(0)^{2}}{1-\pi}\right],\\ &=\mathbb{E}\left[\left(\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+a^{2}_{i}(1)+a^{2}_{i}(0)\right)\right],\\ &\geq\mathbb{E}\left[\left(\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+2\mathbb{E}[a_{i}(1)^{2}]^{1/2}\mathbb{E}[a_{i}(0)^{2}]^{1/2}\right)\right],\\ &=\tilde{\sigma}^{2}_{D},\end{split}$
where the inequality is from the AM-GM inequality. Additionally, to show this is conservative for $\bar{\sigma}^{2}_{D}$
$\begin{split}\bar{\sigma}_{D}^{2}&=\pi\left(1-\pi\right)\mathbb{E}\left[\left(\frac{\varepsilon_{i}(1)}{\pi}+\frac{\varepsilon_{i}(0)}{1-\pi}-\left(\Delta_{i}(1,p^{*}_{\pi})-\Delta_{i}(0,p^{*}_{\pi})\right)\right)^{2}\right],\\ &=\pi\left(1-\pi\right)\mathbb{E}\left[\left(\frac{a_{i}(1)}{\pi}+\frac{a_{i}(0)}{1-\pi}\right)^{2}\right],\\ &=\mathbb{E}\left[\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+2a_{i}(1)a_{i}(0)\right],\\ &\leq\mathbb{E}\left[\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+2\sqrt{\mathbb{E}[a_{i}^{2}(1)]\mathbb{E}[a_{i}^{2}(0)]}\right],\\ &=\tilde{\sigma}^{2}_{D},\end{split}$
where the inequality is from the Cauchy-Schwarz inequality.
D.3 Convergence Rate of $\tau_{\text{AIE}}$ to $\tau^{*}_{\text{AIE}}$
@@
D.4 Observing a Sub-Sample of the Market
@@
D.5 Alternative Estimator of AIE Variance
In analyzing the asymptotic properties of $\sigma^{2}_{I}$ in the proof of Theorem 7 in Appendix B.5, we dropped a term that was asymptotically negligible, which was $\sqrt{n}h_{n}(\xi_{y}\xi_{z}^{-1}(\hat{\tau}^{z}_{\text{ADE}}-\tau^{*,z}_{\text{ADE}})).$
$\tilde{\sigma}^{2}_{I}$ adds a plug-in estimator for the second term, including an estimator for the variance of $\hat{\tau}^{z}_{\text{ADE}}$. $\tilde{\sigma}^{2}_{I}=\hat{\sigma}^{2}_{I}+h^{2}_{n}\hat{\gamma}^{\top}\hat{\sigma}^{2}_{z,D}\hat{\bm{\gamma}},$ where $\hat{\sigma}^{2}_{z,D}=\frac{1}{n}\sum\limits_{i=1}^{n}B_{i}B_{i}^{\prime}$ and the $J\times 1$ vector $\hat{B}_{i}=\frac{W_{i}\hat{\varepsilon}^{z}_{i}(1)}{\hat{\pi}}-\frac{(1-W_{i})\hat{\varepsilon}^{z}_{i}(0)}{1-\hat{\pi}}-[\hat{\xi}_{z1}-\hat{\xi}_{z0}]^{\top}\hat{\xi}_{z}^{-1}{Z}_{i}$. ${\varepsilon}^{z}_{i}(1)=Z_{i}-\frac{1}{n_{w}}\sum\limits_{i:W_{i}=w}Z_{i}$. Last, the $J\times J$ matrix $\hat{\xi}_{zw}$ for $w\in\{0,1\}$ are estimated from regressions of $Z_{i}$ on $U_{i}$ using only observations such that $W_{i}=w$. We found in simulations that this second-order correction leads to better coverage properties at smaller sample sizes.